Find the position vector of the point in space P that lies on the line AB such that |AP|/|PB|= m/n with m, n is a real number. Find the position vector of P if vector A=(1,2,1), B(3,-1,2), m=3 and n=2


1
Expert's answer
2022-05-22T17:59:27-0400

Let "O" be the origin representing the position vectors "\overrightarrow{OA}"  and "\overrightarrow{OB}"  with respect to the two points "A"  and "B" . Let "AB" be divided by a third point "P"  internally in the ratio "m:n."


"\dfrac{AP}{BP}=\dfrac{m}{n}=>n\overrightarrow{AP}=m\overrightarrow{PB}"

"n(\overrightarrow{OP}-\overrightarrow{OA})=m(\overrightarrow{OB}-\overrightarrow{OP})"

"\overrightarrow{OP}=\dfrac{m\overrightarrow{OB}+n\overrightarrow{OA}}{n+m}"

Given


"\overrightarrow{OA}=\langle1,2,1 \rangle, \overrightarrow{OB}=\langle3,-1,2 \rangle"

"m=3, n=2"

"\overrightarrow{OP}=\dfrac{3\langle1,2,1 \rangle+2\langle3,-1,2 \rangle}{2+3}=\langle1.8,0.8,1.4 \rangle"

"\overrightarrow{OP}=\langle1.8,0.8,1.4 \rangle"